OpenStudy (mchilds15):

What is the equation in point-slope form of the line passing through (-2, -5) and (2, 3)? I've narrowed it down to two options but I don't know if i messed up somewhere of what.. Help??

OpenStudy (mchilds15):

It's either y+5=-2(x+2) OR y-3=2(x-2)

OpenStudy (mchilds15):

@Ciarán95

OpenStudy (mchilds15):

Never mind, I got it